# RULSurv: A probabilistic survival-based method for early censoring-aware prediction of remaining useful life in ball bearings

### v2

@@ 5.5 Stratified survival curves
Figure 7 shows the predicted mean survival curve and two curves where the features have been stratified according to their 25th, 50th and 75th percentiles from left to right. The $\mathcal{D}_{C1}$ dataset is used for this experiment with 25% censoring. Recall that the survival function, $S(t)$, is the probability that the event occurs later than time $t$, so higher values are associated with lower predicted risk of event (failure). We see that the predicted survival curves have notable differences based on the feature’s value and that some feature values lead to significantly lower survival probabilities. In fact, across all features evaluated in this plot, the RSF model produces more favorable survival outcomes when the value is below the median and unfavorable outcomes when it is above (center column in Figure 7). For example, the root mean square (RMS) value represents the amount of energy in a signal. As a bearing starts to deteriorate, the RMS value will increase (center row in Figure 7). This happens because the number of peaks increases, which increases the total energy of the signal.
Typically, during the initial phases of a mechanical fault, the RMS value exhibits minimal alteration due to the limited shift in the total energy of the signal. However, as the bearing starts to deteriorate, the RMS value tends to show a more pronounced increase. This way of stratification can provide insight into the learned representation between the inputs and the outputs, and can create counterfactual explanations: If the energy of the signal was less, the bearing would have failed later.
@@ 5.6 Individual RUL prediction
We use the BNNSurv model to make individual RUL predictions per bearing. This is also called the individual survival distribution (ISD). In this case, the ISD is a representation of the survival probability for a single bearing at different time points based on its current features. For each operating condition, we have a total of five bearings, when sampling from the $X$-axis, as suggested by [63]. To predict the ISD of a single bearing, we train $M$ models, where $M$ is the number of operating conditions multiplied by the number of bearings. For example, to predict the ISD of the bearing ($i,j$), where $i$ is the operating condition and $j$ is the bearing index, we train the model $M_{i,j}$ in the dataset $\mathcal{D}_{i}^{-j}$, where the notation $-j$ denotes that the dataset excludes the $j$th bearing. We then predict the ISD by taking a single sample from bearing ($i,j$) at the beginning of its lifetime. We repeat this process iteratively to estimate the ISD for all bearings. Figure 8 shows the predicted ISDs and the upper and lower bounds of the credible intervals (CrIs), estimated by drawing 100 samples from the predictive distribution. We follow [45] and take the median of the survival curve (0.5) as the predicted time to event, $\hat{t}_{i}$, and compare that with the observed event time, $t_{i}$. Although the event detector always detects the event of interest before the end of life (see Table 5), the RUL can be predicted before or after this point. We see a lot of uncertainty in many of the predictions (wide credible intervals), especially for the C1 operating condition due to the low amount of training data available (left column), and there are many irregularities in the predicted RUL, especially for conditions C2 and C2 (center and right columns), which the model cannot reliably predict from just a single sample in this case. Ideally, the predicted time to event should match the actual time to event and preferably before the end of life, so that predictive maintenance can be initiated in time. Depending on the industrial application and the costs of predictive maintenance, overestimating RUL is often worse than underestimating it.
@@ 5.7 Comparison to SOTA
where ActRUL($T_{k}$) is the actual RUL at inspection time $T_{k}$, and RUL($T_{k}$) is the estimated RUL. The closer the CRA is to one, the more accurate the RUL prediction is.
@@
Table 10 presents the performance evaluation results of the six commonly-used RUL prediction methods and the proposed method on the XJTU-SY dataset. Since none of the studied methods provided their source code, it was not possible to reproduce the results, so the CRA is reported as in [63]. For the proposed method, we set the inspection time $T_{k}=1$, i.e., at the beginning of the recording, and take a single sample the $X$-axis. We then train a Random Survival Forests model based on the features and time-to-event information from the remaining four bearings and make the RUL prediction on the holdout bearing, i.e., training a model for each individual bearing iteratively.
@@
The results show that our method gives state-of-the-art predictive performance for operating condition C1. We attribute this to the relatively short event horizon of the bearings in this category; by the time of inspection, the bearings are already close to failure, as indicated by the sensor readings. Consequently, the data collected at this stage are highly predictive of the imminent event. Under a decreased load with a longer event horizon, i.e., conditions C2 and C3, sampling data from the beginning of the recording is no longer as indicative of when the bearing is going to fail. We see that the predictive errors fluctuate a lot and the proposed method generally performs worse than the Hybrid model [63] under conditions C2 and C3. We attribute this behavior to the variance in the XJTU-SY dataset, as some bearings have entirely different lifespans, and also to the lack of good predictive features at inspection time. At $T_{k}=1$, when the bearing is still working fine, the prediction might be less reliable since there is little or no evidence of wear or damage at that early point. On the other hand, sampling from the bearing at a later point in time than $T_{k}=1$ may introduce survivorship bias in this case: if we only make predictions after a certain amount of time has passed (e.g., ten minutes), we are excluding bearings that might have failed earlier, leading to survivorship bias. This can skew the data toward more durable bearings and give an overly optimistic view of how long a specific bearing would last, since we are assuming that the bearing survives past a certain minute mark. In this particular example, we also have fewer training samples at our disposal than the literature methods, since we estimate the RUL as the time to failure rather than the end of life. In some cases, this means that we only have a quarter of the training data available compared to the other methods (see Table 5).
@@
We would have liked to compare our method with [66], who also used the XJTU-SY dataset and proposed a compelling two-stage prediction method; however, the authors only provided aggregated results of the CRA. Similarly, [71] proposed a multiscale encoder-decoder consisting of a CNN and a multi-head attention mechanism, but the authors only provided a single RMS error, not an individual error for each bearing. [58] proposed a deep adaptive transformer model, which supposedly provides better forecasting performance and solves the problem of vanishing gradients in existing RNN architectures, but the reported relative accuracy is only for two of the fifteen bearings in the XJTU-SY dataset. Also, the authors did not provide source code to reproduce their results.
@@ 6.1 Limitations
We propose an event detection algorithm based on the KL divergence. This offers interpretability, as opposed to using a DL architecture, but also requires manual feature engineering and fine-tuning, since we have to carefully select the frequency bands and calibrate the threshold function to avoid false positives and false negatives. In addition, there are some challenges in using the proposed frequency spectra to identify anomalies:
@@
    In all rotating machinery, the observed vibration signature will be a sum of contributions from mechanical components in the system; bearing defects, bearing looseness, misalignment of shafts from e.g., the motor to the gearbox, imbalances in the rotating parts or gear-teeth meshing. These contributions will interact, and the resultant characteristic frequencies can add or subtract with a magnitude determined by the sensor location and how mobile each component is, obfuscating the informative frequencies.
@@
    Additional factors, such as the quality of lubrication, the roughness of the rolling element surface, and changes in surface conformance, do not manifest themselves as a characteristic cyclic frequency, which makes them difficult to detect with traditional model-based spectral analysis or classical data-driven ML methods.
@@
Our method supports several survival models (CoxPH, GBSA, and BNNSurv) that assume proportional hazards, i.e., the relative risk of an event is independent of time. This assumption is easier to satisfy when the event horizon is short. In an industrial setting, where bearings usually last longer than under artificially accelerated degradation, the event time distribution may have temporal dependencies at different times that may not be easily captured by a proportional hazard model.
Based on the estimated survival curve, we use the median of the curve (0.5) as the predicted time to event. Other approaches have been proposed, such as using the area under the survival function (AUC) as the predicted time, or simulating the survival times based on some probability distribution [4]. Although it is unlikely that a single bearing will fail at the exact time the survival curve reaches 0.5, a population of bearings will fail at that time on average, given a sufficiently large sample size and that the survival curves are D-calibrated. The XJTU-SY dataset only comes with five bearings per operating condition and there is a pronounced variability between sensor readings, too, even for bearings that are tested under the same load configuration. This leads to inaccuracies when predicting the exact RUL.
@@ 6.2 Industrial applicability
Moreover, the effectiveness of the proposed method is based on the assumption that defects progress in a predictable and gradual manner, starting from the initial stages and advancing through a series of deterioration steps (see Fig. 1). This steady development of bearing defects, whether they occur alone or in combination, would facilitate detailed monitoring and data collection in industrial settings, due to the gradual changes in vibration patterns. However, despite this gradual progression, the variation in vibration signatures among even similar rotating machinery can be significant, as discussed in Section 6.1. In contrast, the vibration patterns observed in the XJTU-SY dataset are the result of excessive load on the bearings, preventing proper lubrication and leading to immediate surface damage and swift degradation of the raceway. Although the proposed method performed well on the XJTU-SY dataset, which features bearings that undergo accelerated degradation, industrial settings often present a different challenge, where degradation is less apparent and wear develops gradually over time. Given this disparity, practitioners need to carefully select the machine-specific thresholds for the event detector and have a sufficient hold-out of set of multiple bearings to evaluate the performance of the predictive model. Whenever the model is used for a novel application, the event detector and the predictive model will have to be reevaluated. In summary, the extension to other types of data sets and industrial applications poses the following challenges:
@@
    Industrial sensor data is often noisier and less reliable than controlled experimental data. In an industrial setting, sensor failures, environmental factors, and other variables could introduce noise into the training dataset. In the presented work, we do not perform any outlier detection before training the models. However, if significant noise or anomalies are present in the input data, incorporating an anomaly detection algorithm may be necessary.
@@
    The survival models are only evaluated in the XJTU-SY dataset, where bearings are subject to accelerated degradation. In a new dataset, models will have been retrained and reevaluated in a hold-out set. Although, using either the MTLR or the BNNSurv model, the learned parameters (weights and biases) can be used as initial weights when training on a new dataset (i.e., pretraining).
@@ 6.3 Scalability
In this study, we have proposed several survival models for RUL prediction, both neural networks (MTLR and BNNSurv) and non-neural networks (CoxPH, GBSA, and RSF). In terms of model scalability to larger datasets with more observations and a larger number of features, the scalability of the non-neural networks is limited by memory constraints. As datasets grow larger, these methods become less practical as they have to process the entire dataset for each gradient update during training. On the other hand, a neural network only loads small batches of the entire training set into memory during training, allowing the system to handle larger datasets efficiently without requiring a proportional increase in memory. This makes MTLR and BNNSurv suitable for training on large-scale datasets that exceed the available memory on a given system. Moreover, using minibatches also allows for better hardware utilization, such as GPUs, as they can process multiple samples in parallel.
@@ 7 Conclusion
We have introduced a novel method for event detection and RUL estimation in ball bearings that naturally supports censored data. The key outcomes of the proposed method are as follows: (1) It offers an interpretable event detection algorithm based on Kullback-Leibler (KL) divergence to annotate a ball bearing dataset with the time a bearing starts to fail, before the end of life. (2) It provides a flexible set of survival models to predict the RUL, which is guaranteed to decrease monotonically as a function of time. (3) It uses the median of the survival curve as the predicted time to event, providing a quantitative and probabilistic estimate of the event occurrence. (4) It gives state-of-the-art predictive performance on the XJTU-SY dataset under the highest load condition. (5) It can make use of uncertainty-aware survival methods, e.g., Bayesian methods, which can predict an individual survival distribution with credible intervals for each bearing. The findings of this work have significant implications for real-life predictive maintenance and reliability engineering: (1) Given a ball bearing dataset with fully-observed recordings, the proposed event detector can label each bearing with its time to failure (i.e., where the bearing starts to malfunction), not its end of life. Thus, any machine learning model trained on this dataset will learn to predict the time to failure instead of the end of life. (2) By leveraging censored data in predictive modeling, the time-to-event estimates are evidently more accurate than if we were to use only data from bearings with a known failure time. (3) The proposed method includes five survival models, suitable for modeling under different conditions e.g., under nonproportional hazards or in nonlinear risk applications. In general, these three contributions can lead to better RUL predictions, which in turn can help industries such as manufacturing, transportation, and energy reduce downtime, lower maintenance costs, and improve operational efficiency.

### v3

@@ 5.5 Stratified survival curves
Figure 7 shows the predicted mean survival curve and two curves where the features have been stratified according to their 25th, 50th and 75th percentiles from left to right. The $\mathcal{D}_{C1}$ dataset is used for this experiment with 25% censoring. Recall that the survival function, $S(t)$, is the probability that the event occurs later than time $t$, so higher values are associated with a lower predicted risk of the event (failure). We see that the predicted survival curves have notable differences based on the feature’s value and that some feature values lead to significantly lower survival probabilities. In fact, across all features evaluated in this plot, the RSF model produces more favorable survival outcomes when the value is below the median and unfavorable outcomes when it is above (center column in Figure 7). For example, the root mean square (RMS) value represents the amount of energy in a signal. As the bearing begins to deteriorate, the RMS value will increase (center row in Figure 7). This happens because the number of peaks increases, which increases the total energy of the signal.
Typically, during the initial phases of a mechanical fault, the RMS value exhibits minimal alteration due to the limited shift in the total energy of the signal. However, as the bearing begins to deteriorate, the RMS value tends to show a more pronounced increase. This way of stratification can provide insight into the learned representation between the inputs and the outputs, and can create counterfactual explanations: If the signal energy was less, the bearing would have failed later.
@@ 5.6 Individual RUL prediction
We use the BNNSurv model to make individual RUL predictions per bearing. This is also called the individual survival distribution (ISD). In this case, the ISD is a representation of the survival probability for a single bearing at different time points based on its current features. For each operating condition, we have a total of five bearings, when sampling from the $X$-axis, as suggested by [64]. To predict the ISD of a single bearing, we train $M$ models, where $M$ is the number of operating conditions multiplied by the number of bearings. For example, to predict the ISD of the bearing ($i,j$), where $i$ is the operating condition and $j$ is the bearing index, we train the model $M_{i,j}$ in the dataset $\mathcal{D}_{i}^{-j}$, where the notation $-j$ denotes that the dataset excludes the $j$th bearing. We then predict the ISD by taking a single sample from bearing ($i,j$) at the beginning of its lifetime. We repeat this process iteratively to estimate the ISD for all bearings. Figure 8 shows the predicted ISDs and the upper and lower bounds of the credible intervals (CrIs), estimated by drawing 100 samples from the predictive distribution. We follow [47] and take the median of the survival curve (0.5) as the predicted time-to-event, $\hat{t}_{i}$, and compare it with the observed event time, $t_{i}$. Although the event detector always detects the event of interest before the end of life (see Table 5), the RUL can be predicted before or after this point. We see a lot of uncertainty in many of the predictions (wide credible intervals), especially for the C1 operating condition due to the low amount of training data available (left column), and there are many irregularities in the predicted RUL, especially for conditions C2 and C2 (center and right columns), which the model cannot reliably predict from just a single sample in this case. Ideally, the predicted time-to-event should match the observed time-to-event and preferably before the end of life, so that predictive maintenance can be initiated in time. Depending on the industrial application and the costs of predictive maintenance, overestimating RUL is often worse than underestimating it.
@@ 5.7 Comparison to SOTA
where ActRUL($T_{k}$) is the actual RUL at inspection time $T_{k}$, and RUL($T_{k}$) is the estimated RUL. The closer the CRA to one, the more accurate the prediction.
@@
Table 10 presents the performance evaluation results of the six commonly-used RUL prediction methods and the proposed method on the XJTU-SY dataset. Since none of the methods studied provided their source code, it was not possible to reproduce the results, so the CRA is reported as in [64]. For the proposed method, we set the inspection time $T_{k}=1$, i.e., at the beginning of the recording, and take a single sample the $X$-axis. We then train a Random Survival Forests model based on the features and time-to-event information of the remaining four bearings and make the RUL prediction on the holdout bearing, i.e., training a model for each individual bearing iteratively.
@@
The results show that our method achieves state-of-the-art predictive performance under operating condition C1. We attribute this to the relatively short event horizon of the bearings in this category. By the time of inspection, the bearings are already nearing failure, as evidenced by the sensor readings. As a result, the data collected at this stage are highly indicative of the impending event. Under a decreased load with a longer event horizon, i.e., conditions C2 and C3, sampling data from the beginning of the recording is no longer indicative of when the bearing is going to fail. We see that the predictive errors fluctuate a lot and the proposed method generally performs worse than the Hybrid model [64] under conditions C2 and C3. We attribute this behavior to the variance in the XJTU-SY dataset, as some bearings have entirely different lifespans, and also to the lack of good predictive features at inspection time. At $T_{k}=1$, when the bearing is still working properly, the prediction might be less reliable since there is little or no evidence of wear or damage at that early point. On the other hand, sampling from the bearing at a later point in time than $T_{k}=1$ may introduce survivorship bias in this case: If we only make predictions after a certain amount of time has passed (e.g., ten minutes), we exclude bearings that might have failed earlier, leading to survivorship bias. This can skew the data toward more durable bearings and give an overly optimistic view of how long a specific bearing would last, since we are assuming that the bearing survives past a certain minute mark. In this particular example, we also have fewer training samples at our disposal than the literature methods, since we estimate the RUL as the time to failure rather than the end of life. In some cases, this means that we only have a quarter of the training data available compared to the other methods (see Table 5).
@@
We would have liked to compare our method with [67], who also used the XJTU-SY dataset and proposed a compelling two-stage prediction method; however, the authors only provided aggregated results of the CRA. Similarly, [72] proposed a multiscale encoder-decoder consisting of a CNN and a multi-head attention mechanism, but the authors only provided a single RMS error, not an individual error for each bearing. [59] proposed a deep adaptive transformer model, which supposedly provides better forecasting performance and solves the problem of vanishing gradients in existing RNN architectures, but the reported relative accuracy is only for 2 of the 15 bearings in the XJTU-SY dataset. In addition, the authors did not provide source code to reproduce their results.
@@ 6.1 Limitations
We propose an event detection algorithm based on the KL divergence. This offers interpretability, as opposed to using a DL architecture, but also requires manual feature engineering and fine-tuning, since we have to carefully select the frequency bands and calibrate the threshold function to avoid false positives and false negatives. In addition, there are some challenges in using the proposed frequency spectra to identify anomalies.
@@
    In all rotating machinery, the observed vibration signature will be a sum of contributions from mechanical components in the system; bearing defects, bearing looseness, misalignment of shafts from e.g., the motor to the gearbox, imbalances in the rotating parts, or gear-teeth meshing. These contributions will interact and the resultant characteristic frequencies can be added or subtracted with a magnitude determined by the sensor location and how mobile each component is, obfuscating the informative frequencies.
@@
    Additional factors, such as the quality of lubrication, the roughness of the surface of the rolling element, and changes in surface conformance, do not manifest themselves as characteristic cyclic frequencies, making them difficult to detect with traditional model-based spectral analysis or classical data-driven ML methods.
@@
Our method supports several survival models (CoxPH, GBSA, and BNNSurv) that assume proportional hazards, i.e., the relative risk of an event is independent of time. This assumption is easier to satisfy when the event horizon is short. In an industrial setting, where bearings usually last longer than under artificially accelerated degradation, the time-to-event distribution may have temporal dependencies at different times that may not be easily captured by a proportional hazard model.
Based on the estimated survival curve, we use the median of the curve (0.5) as the predicted time-to-event. Other approaches have been proposed, such as using the area under the survival function (AUC) as the predicted time, or simulating the survival times based on some probability distribution [4]. Although it is unlikely that a single bearing will fail at the exact time the survival curve reaches 0.5, a population of bearings will fail at that time on average, given a sufficiently large sample size and that the survival curves are D-calibrated. The XJTU-SY dataset includes only five bearings per operating condition, and sensor readings show considerable variability even among bearings tested under the same load configuration. This variability contributes to inaccuracies in predicting the exact RUL.
@@ 6.2 Industrial applicability
Moreover, the effectiveness of the proposed method is based on the assumption that defects progress in a predictable and gradual manner, starting from the initial stages and advancing through a series of deterioration steps (see Fig. 1). This steady development of bearing defects, whether they occur alone or in combination, would facilitate detailed monitoring and data collection in industrial settings, due to the gradual changes in vibration patterns. However, despite this gradual progression, the variation in vibration signatures among even similar rotating machinery can be significant, as discussed in Section 6.1. In contrast, the vibration patterns observed in the XJTU-SY dataset are the result of excessive load on the bearings, preventing proper lubrication and leading to immediate surface damage and rapid degradation of the raceway. Although the proposed method performed well on the XJTU-SY dataset, which features bearings that undergo accelerated degradation, industrial settings often present a different challenge, where degradation is less apparent and wear develops gradually over time. Given this disparity, practitioners need to carefully select the machine-specific thresholds for the event detector and have a sufficient hold-out of set of multiple bearings to evaluate the performance of the predictive model. Whenever the model is used for a novel application, the event detector and the predictive model will have to be reevaluated. In summary, the extension to other types of data sets and industrial applications poses the following challenges:
@@
    Industrial sensor data are often noisier and less reliable than controlled experimental data. In an industrial setting, sensor failures, environmental factors, and other variables could introduce noise into the training dataset. In the presented work, we do not perform any outlier detection before training the models. However, if significant noise or anomalies are present in the input data, incorporating an anomaly detection algorithm may be necessary.
@@
    The survival models are only evaluated in the XJTU-SY dataset, where bearings are subject to accelerated degradation. In a new dataset, models will have been retrained and reevaluated in a hold-out set. However, using either the MTLR or the BNNSurv model, the learned parameters (weights and biases) can be used as initial weights when training on a new dataset (i.e., pretraining).
@@ 6.3 Scalability
In this study, we have proposed several survival models for RUL prediction, both neural networks (MTLR and BNNSurv) and non-neural networks (CoxPH, GBSA, and RSF). In terms of model scalability to larger datasets with more observations and a larger number of features, the scalability of the non-neural networks is limited by memory constraints. As datasets grow larger, these methods become less practical as they have to process the entire dataset for each gradient update during training. However, a neural network only loads small batches of the entire training set into memory during training, allowing the system to handle larger datasets efficiently without requiring a proportional increase in memory. This makes MTLR and BNNSurv suitable for training on large-scale datasets that exceed the available memory on a given system. Moreover, using minibatches also allows for better hardware utilization, such as GPUs, as they can process multiple samples in parallel.
@@ 7 Conclusion
We have introduced a novel method for event detection and RUL estimation in ball bearings that naturally supports censored data. The key outcomes of the proposed method are as follows: (1) It offers an interpretable event detection algorithm based on Kullback-Leibler (KL) divergence to annotate a ball bearing dataset with the time a bearing starts to fail, before the end of life. (2) It provides a flexible set of survival models to predict the RUL, which is guaranteed to decrease monotonically as a function of time. (3) It uses the median of the survival curve as the predicted time-to-event, providing a quantitative and probabilistic estimate of the event occurrence. (4) It gives state-of-the-art predictive performance on the XJTU-SY dataset under the highest load condition. (5) It can make use of uncertainty-aware survival methods, e.g., Bayesian methods, which can predict an individual survival distribution with credible intervals for each bearing. The findings of this work have significant implications for real-life predictive maintenance and reliability engineering: (1) Given a ball bearing dataset with fully-observed recordings, the proposed event detector can label each bearing with its time to failure (i.e., where the bearing starts to malfunction), not its end of life. Thus, any machine learning model trained on this dataset will learn to predict the time to failure instead of the end of life. (2) By leveraging censored data in predictive modeling, the time-to-event estimates are evidently more accurate than if we were to use only data from bearings with a known failure time. (3) The proposed method includes five survival models, suitable for modeling under different conditions e.g., under nonproportional hazards or in nonlinear risk applications. In general, these three contributions can lead to better RUL predictions, which in turn can help industries such as manufacturing, transportation, and energy reduce downtime, lower maintenance costs, and improve operational efficiency.